# Causal Deciphering and Inpainting in Spatio-Temporal Dynamics via Diffusion Model

### v1

@@ Abstract
Spatio-temporal (ST) prediction has garnered a De facto attention in earth sciences, such as meteorological prediction, human mobility perception. However, the scarcity of data coupled with the high expenses involved in sensor deployment results in notable data imbalances. Furthermore, models that are excessively customized and devoid of causal connections further undermine the generalizability and interpretability. To this end, we establish a causal framework for ST predictions, termed CaPaint, which targets to identify causal regions in data and endow model with causal reasoning ability in a two-stage process. Going beyond this process, we utilize the front door adjustment to specifically address the sub-regions identified as non-causal in the upstream phase. Specifically, we employ a novel image inpainting technique. By using a fine-tuned unconditional Diffusion Probabilistic Model (DDPM) as the generative prior, we in-fill the masks defined as environmental parts, offering the possibility of reliable extrapolation for potential data distributions. CaPaint overcomes the high complexity dilemma of optimal ST causal discovery models by reducing the data generation complexity from exponential to quasi-linear levels. Extensive experiments conducted on five real-world ST benchmarks demonstrate that integrating the CaPaint concept allows models to achieve improvements ranging from 4.3$\sim$77.3%. Moreover, compared to traditional mainstream ST augmenters, CaPaint underscores the potential of diffusion models in ST enhancement, offering a novel paradigm for this field. Our project is available at https://anonymous.4open.science/r/12345-DFCC.
@@ 1 Introduction
In light of this, we propose a general causal structure plugin, termed CaPaint, designed to decipher causal regions in ST data without adding extra computational cost, while intervening in non-causal areas to boost the model’s generalizability and interpretability. Specifically, our method employs a straightforward approach to causal discovery, utilizing a vision transformer architecture [32] for self-supervised ST data reconstruction. During reconstruction, we leverage attention scores from the self-attention mechanism [22] to map onto important causal patches, thus endowing the model with interpretability. By ranking the entire set of importance scores, we define those with lower scores as environmental patches, which contribute minimally to the model. Building on this, we perform causal interventions in these environmental areas to aid the model in understanding more latent, complex, and imperceptible distributions, thereby enhancing the overall generalizability of the model (see Figure 1). Concretely, we mask trivial patches and perform denoising process via Denoising Diffusion Probabilistic Models (DDPM) [23, 31], this also can be interpreted as an image inpainting approach.
@@
  • We conduct extensive experiments across five diverse and representative datasets from different domains, utilizing seven backbone models to assess the effectiveness of the CaPaint method. The empirical results demonstrate that CaPaint consistently enhances performance on all tested datasets and across all backbone models (4.3$\sim$77.3%).
@@ 2 Related work & Technical Background
Spatio-temporal Predictive Learning: Various architectures have achieved significant predictive performance in ST domain, which can primarily be categorized as follows: CNN-based models utilize convolutional layers to effectively capture spatial features [44, 47, 73, 9]. RNN-based models, are capable of processing temporal sequence data and are well-suited for understanding temporal changes, showing excellent performance in the prediction of action continuity [65, 76, 82, 68]. GNN-based models effectively capture spatial dependencies and temporal dynamics in data, making them suitable for complex tasks involving geographic locations and temporal changes [43, 24, 35, 16, 95, 94, 79]. Transformer-based models employ self-attention mechanisms to process sequential data in parallel, enhancing the learning of long-term dependencies, and have been used for ST data prediction in complex scenarios [2, 18, 85, 87, 10, 84].
@@
Generative models especially diffusion-based model has gained significant popularity particularly in image and video generation [23]. Sampling optimization algorithms have been used to accelerate the sampling process of diffusion models, significantly reducing the number of steps while improving efficiency. [63, 40]. Additionally, generative models have also been applied to 3D scene generation and point cloud processing, as demonstrated in [39, 29, 69, 70, 21]
@@ 3.1 Causal Deciphering
To address issues of ST data scarcity and poor transferability, we examine the evaluation process using a Structural Causal Model (SCM) [51], as shown in Fig 3. We represent abstract data variables by nodes, with directed links symbolizing causality. The SCM illustrates the interaction among variables through a graphical definition of causation, demonstrating the interconnected nature of these elements. As depicted in the left part, existing SOTA causal model, NuwaDynamics, employs the backdoor adjustment to enhance the model’s generalization performance:
@@
In general, a model ${{\mathcal{F}}_{\emptyset}}$ trained using Empirical Risk Minimization (ERM) often struggles to generalize to the test data $\mathcal{D}_{te}\sim\mathbb{P}_{te}$. Such distribution shifts are often induced by variations in environmental patches. Hence, addressing the confounding effect caused by the environmental confounder is crucial. NuwaDynamics employs backdoor adjustment techniques to perturb the environmental components. Concretely, NuwaDynamics enhances the model’s potential to observe a broader range of latent distributions by forcibly perturbing the environmental variables ${{\mathcal{X}}_{\backslash{\mathcal{C}}}}$ (called do-calculus [50] operator). Unfortunately, traversing all environmental variables is quite challenging. Although NuwaDynamics uses Gaussian sampling to mitigate the issue of complexity, controlling Gaussian sampling in temporal sequence operations is particularly difficult. It requires meticulous adjustment of mean and variance to ensure a balance between the number of environmental samples and the training burden. Worse still, by traversing all environments, it likely violates underlying properties, including distribution shift content and nonexistent scenarios [90]. To address this issue, we employ front-door adjustment, as illustrated in the right half of the Fig 3:
@@ 3.4 ST Sequence Sampling Modeling
Previous work [78] assumed that the closer the time point is to the present, the greater its influence, and thus used Gaussian sampling to select more ST data closer to the current time point. However, we argue that uniform sampling can better enhance the model’s generalization ability. To introduce a trade-off between computational efficiency and accuracy, we implement a strategy inspired by the dropout technique commonly used in neural networks. Specifically, we employ a selective dropout mechanism at each time point to randomly exclude certain augmented images from processing. This approach effectively reduces the computational complexity without substantially compromising the model’s performance. The selective dropout process can be mathematically represented as follows:
$\tilde{X}_{t}$ =$=\{x_{t}^{i}\in X_{t}:z_{t}^{i}<p\},\quad p$ $=\min\left(1,\frac{\gamma}{\|X_{t}\|}\right)$ (10)
where $z_{i}^{t}$ is a random variable drawn independently from a uniform distribution $U(0,1)$, and $p$ is the probability of retaining any given image, balancing computational efficiency and model accuracy. The parameter $\gamma$ dynamically adjusts $p$ based on the system’s load to optimize resource utilization.

### v2

@@ Abstract
Spatio-temporal (ST) prediction has garnered a De facto attention in earth sciences, such as meteorological prediction, human mobility perception. However, the scarcity of data coupled with the high expenses involved in sensor deployment results in notable data imbalances. Furthermore, models that are excessively customized and devoid of causal connections further undermine the generalizability and interpretability. To this end, we establish a framework for ST predictions from a causal perspective, termed CaPaint, which targets to identify causal regions in data and endow model with causal reasoning ability in a two-stage process. Going beyond this process, we build on the front door adjustment as the theoretical foundation to specifically address the sub-regions identified as non-causal in the upstream phase. By using a fine-tuned unconditional Diffusion Probabilistic Model (DDPM) as the generative prior, we in-fill the masks defined as environmental parts, offering the possibility of reliable extrapolation for potential data distributions. CaPaint overcomes the high complexity dilemma of optimal ST causal discovery models by reducing the data generation complexity from exponential to quasi-linear levels. Extensive experiments conducted on five real-world ST benchmarks demonstrate that integrating the CaPaint concept allows models to achieve improvements ranging from 3.7$\sim$77.3%. Moreover, compared to traditional mainstream ST augmenters, CaPaint underscores the potential of diffusion models in ST data augmentation, offering a novel paradigm for this field. Our project is available at CaPaint.
@@ 1 Introduction
In light of this, we propose a general causal structure plugin, termed CaPaint, designed to decipher causal regions in ST data without adding extra computational cost, while intervening in non-causal areas to boost the model’s generalizability and interpretability. Specifically, our method employs a straightforward approach to causal deciphering, utilizing a vision transformer architecture [33] for self-supervised ST data reconstruction. During reconstruction, we leverage attention scores from the self-attention mechanism [23] to map onto important causal patches, thus endowing the model with interpretability. By ranking the entire set of importance scores, we define those with lower scores as environmental patches, which contribute minimally to the model. Building on this, we perform causal interventions in these environmental areas to aid the model in understanding more latent, complex, and imperceptible distributions, thereby enhancing the overall generalizability of the model (see Figure 1). Concretely, we mask trivial regions and perform generation using DDPM [24, 32] fine-tuned on specific ST data, which can also be interpreted as a ST data inpainting approach.
@@
  • We conduct extensive experiments across five diverse and representative datasets from different domains, utilizing seven backbone models to assess the effectiveness of the CaPaint method. The empirical results demonstrate that CaPaint consistently enhances performance on all tested datasets and across all backbone models (3.7$\sim$77.3%).
@@ 2 Related work & Technical Background
Spatio-temporal Predictive Learning: Various architectures have achieved significant predictive performance in ST domain, which can primarily be categorized as follows: CNN-based models utilize convolutional layers to effectively capture spatial features [45, 48, 77, 9]. RNN-based models, are capable of processing temporal sequence data and are well-suited for understanding temporal changes, showing excellent performance in the prediction of action continuity [69, 80, 86, 72]. GNN-based models effectively capture spatial dependencies and temporal dynamics in data, making them suitable for complex tasks involving geographic locations and temporal changes [44, 25, 36, 17, 99, 98, 83, 14]. Transformer-based models employ self-attention mechanisms to process sequential data in parallel, enhancing the learning of long-term dependencies, and have been used for ST data prediction in complex scenarios [2, 19, 89, 91, 10, 88].
@@
Generative models especially diffusion-based model has gained significant popularity particularly in image and video generation [24, 64, 62]. Sampling optimization algorithms have been used to accelerate the sampling process of diffusion models, significantly reducing the number of steps while improving efficiency. [67, 41]. Additionally, generative models have also been applied to 3D scene generation and point cloud processing, as demonstrated in [40, 30, 73, 74, 22, 63]
@@ 3.1 Causal Deciphering
To address issues of ST data scarcity and poor transferability, we examine the evaluation process using a Structural Causal Model (SCM) [52], as shown in Fig 3. We represent abstract data variables by nodes, with directed links symbolizing causality. The SCM illustrates the interaction among variables through a graphical definition of causation, demonstrating the interconnected nature of these elements. As depicted in the left part, NuwaDynamics employs the backdoor adjustment to enhance the model’s generalization performance:
@@
In general, a model ${{\mathcal{F}}_{\emptyset}}$ trained using Empirical Risk Minimization (ERM) often struggles to generalize to the test data $\mathcal{D}_{te}\sim\mathbb{P}_{te}$. Such distribution shifts are often induced by variations in environmental patches. Hence, addressing the confounding effect caused by the environmental confounder is crucial. Backdoor adjustment techniques are employed to perturb the environmental components, thereby enhancing the model’s potential to observe a broader range of latent distributions by forcibly perturbing the environmental variables ${{\mathcal{X}}_{\backslash{\mathcal{C}}}}$ (referred to as the do-calculus [51] operator). Unfortunately, traversing all environmental variables is quite challenging. Although NuwaDynamics uses Gaussian sampling to mitigate the issue of complexity, controlling Gaussian sampling in temporal sequence operations is particularly difficult. It requires meticulous adjustment of mean and variance to ensure a balance between the number of environmental samples and the training burden. Worse still, by traversing all environments, it likely violates underlying properties, including distribution shift content and nonexistent scenarios [94]. To address this issue, we employ front-door adjustment, as illustrated in the right half of the Fig 3:
@@ 3.4 ST Sequence Sampling Modeling
Previous work [82] assumed that the closer the time point is to the present, the greater its influence, and thus used Gaussian sampling to select more ST data closer to the current time point. However, we argue that uniform sampling can better enhance the model’s generalization ability. To enhance computational efficiency while ensuring prediction accuracy, we employ a ST sequence modeling approach that samples at each time point with a fixed probability controlled by the hyperparameter $p$. This method allows us to sample from both original and generated data at each time point, thereby creating a new spatiotemporal sequence. We use two hyperparameters: $p$, which controls the sampling probability, and $r$, which determines the number of generated spatiotemporal sequences, achieving an optimal balance between computational efficiency and prediction accuracy. The specific sampling process can be represented by the following equation:
$X^{\prime}_{t}=\text{Sample}(X_{t},p,r)$ (10)
where $X_{t}$ represents the collection of original and generated data at time point $t$, and $\text{Sample}(X_{t},p)$ denotes the dataset obtained by sampling from $X_{t}$ with probability $p$. The hyperparameter $p$ is directly set as the sampling probability, while $r$ is used to specify the number of generated ST sequences.
@@ 5 Conclusion & Future Work
6 Acknowledgement
This work was supported by National Natural Science Foundation of China (62476224).
@@ Appendix I Visualizations on KTH
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In this paper, we introduce the spatio-temporal causal concept in the data mining realm, aimed at enhancing the reliability and accuracy of financial spatio-temporal prediction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In this work, we systematically discuss the limitations of our research and outline directions for future work.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not include experimental results related to theoretical aspects.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide the code necessary for replicating the studies described in this paper via an anonymous link, and we detail the experimental setup for the replication in the article itself.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: For the datasets disclosed in the article, we have provided information regarding their sources and origins.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: we have specified all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In this paper, we have reported error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In this paper, we provide detailed information about the experimental resources, including GPU configurations used in our studies.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The study presented in this paper conforms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We have provided the societal impacts of the work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper does not address issues related to this aspect.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All creators and original owners of the assets used in our paper, such as code, data, and models, have been properly credited. We have explicitly mentioned the licenses and terms of use for each asset and have ensured full compliance with these terms throughout our research.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The research presented in this paper is not concerned with new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve experiments or research related to human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not address potential risks incurred by study participants.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.